I have to find the solution and graph this I am so confused.
5x+7y=b

Answer by josmiceli(19441)

I think you're confused because you've got a letter where you expected a number to be.
Just get used to it- algebra problems are full of theings like this.
You want to put it in the form y = mx +b. Now this "b" has nothing to do with the b in the problem, let's change y = mx + b to y = mx + k.
5x + 7y = b
subtract 5x from both sides
7y = -5x + b
divide both sides by 7
y = (-5/7)x + (b/7)
now it's in the form y = mx + k. This form helps with graphing, because m is always the slope.
so, (-5/7) is the slope
The thing about negative slopes, the line goes from your upper left down to
your lower right on the graph
just find these two points, and your have the graph
(x, y) = (0, y) this is called the y-intercept
(x, y) = (x, 0) this is called the x-intercept
If you make x = 0 in the equation, you get the first point
It's (0, b/7)
If you make y = 0 in the equation, you get the second point
It's (b/5, 0)
The only solid data you have is slope = -5/7
The graph is ALL the lines with that slope, depending on what you want to pick for b. For instance, if b = 35, then your intercepts are:
(0, 5) and
(7, 0)
The slope never changes, though
hope this helps